# Exact inference via quasi-conjugacy in two-parameter Poisson–Dirichlet hidden Markov models

### v1

@@ 5 Concluding remarks
We developed an exact and tractable inferential framework for nonparametric hidden Markov models where the latent process is a two-parameter Poisson–Dirichlet diffusion and the observations are unlabelled partitions evolving over discrete time. By exploiting Markov duality and combinatorial properties of partition coagulations, we derived recursive updates for both posterior and predictive distributions. These results enabled exact filtering, smoothing, and forecasting algorithms, as well as marginal likelihood-based parameter estimation.
Our approach demonstrates practical scalability and statistical fidelity in modeling time-varying unlabelled partition data. Unlike conventional particle filters, which require simulating high-dimensional latent dynamics and evaluating intractable likelihoods, our method sidesteps both difficulties through structural duality and quasi-conjugacy. The resulting algorithms are not only substantially faster, educing runtime from approximately one hour to under two minutes, but also more stable, yielding coherent posterior trajectories and sharper uncertainty quantification. As shown through synthetic and real-world experiments, the proposed methodology enables tractable, interpretable, and provably exact inference in settings where the latent structure is infinite-dimensional and only partially observed. These properties make it particularly well suited for applications with computational constraints or unobserved individual-level labels, such as social or genetic data.
Appendix A Appendix
@@ A.1 Two-parameter Poisson–Dirichlet diffusions
with domain given by the subalgebra of the space $C(\overline{\nabla})$ of continuous function on $\overline{\nabla}$ generated by $\{\varphi_{m}(x),m\in\mathbb{N}\}$, with $\varphi_{1}(x)\equiv 1$ and $\varphi_{m}(x)=\sum_{i=1}^{\infty}x_{i}^{m}$ for $m\geq 2$. See, e.g., feng2010poisson, Section 5.5. The transition function of the diffusion was found in feng2011functional through a spectral expansion, and admits the representation as infinite mixture of conditional two-parameter Poisson–Dirichlet distributions
@@
where $a_{(m)}=a(a+1)\cdots(a+m+1)$ is the ascending factorial or Pochhammer symbol, with $a_{(0)}:=1$. These $d_{n}(t)$ are the transition probabilities of a death process on $\mathbb{Z}_{+}=\{0,1,\dots\}$ that has entrance boundary at $\infty$ and infinitesimal transition rates $n(\theta+n-1)/2$ from $n$ to n1n-1. This process is in fact the block-counting process associated to Kingman’s coalescent with mutation (tavare1984line).
@@
at all $t>0$, so that $\overline{\nabla}\setminus\nabla$ acts as an entrance boundary. Then, if $X(0)\in\nabla$ almost surely, we can take $\nabla$ as the process state space. For this reason we assumed $X(0)\sim\mathrm{PD}_{\alpha,\theta}$, which implies $X(0)\in\nabla$ with probability one (pitman2006combinatorial).
@@ A.2.1 Two-parameter Poisson–Dirichlet partition structures
An integer partition of $n\in\mathbb{N}$ is a nonincreasing vector $\pi=(\pi_{1},\ldots,\pi_{l(\pi)})$ of positive integers with $|\pi|:=\sum_{i=1}^{l(\pi)}\pi_{i}=n$. See Figure 2 for examples of Young diagram representation of partitions. Let $\mathcal{P}_{n}$ be the set of partitions of $n$ and $\mathcal{P}:=\bigcup_{n\geq 1}\mathcal{P}_{n}$. For $\pi\in\mathcal{P}$, write $a(\pi):=(a_{j}(\pi),\,j\geq 1)$ where
$a_{j}(\pi):=\sum_{i=1}^{l(\pi)}\mathbbm{1}\{\pi_{i}=j\}$ (27)
@@
A partition structure (kingman_representation_1978) is a family $\{P_{n}:n\in\mathbb{N}\}$ with $P_{n}$ a probability on $\mathcal{P}_{n}$ such that $P_{n}$ is the marginal obtained by removing uniformly at random one square from the Young diagram of $\Pi\sim P_{n+1}$.
@@
$C(\pi)\;=\;\binom{|\pi|}{\pi_{1},\ldots,\pi_{l(\pi)}}\;\frac{1}{a_{1}(\pi)!\cdots a_{|\pi|}(\pi)!},$ (28)
and we use the rising factorial notation $q_{(b)}:=q(q+1)\cdots(q+b-1)$ (with the convention $q_{(0)}:=1$). A random partition $\Pi$ of size $n$ with pmf $\{\mathrm{PSF}_{\alpha,\theta}(\pi),\ \pi\in\mathcal{P}_{n}\}$ is denoted $\Pi\sim\mathrm{CRP}(n)$. Equivalently, $\Pi$ can be generated by the $(\alpha,\theta)$ Chinese restaurant process (CRP): the $i$th item joins an existing block of current size $n_{j}$ with probability $(n_{j}-\alpha)/(\theta+i-1)$, or starts a new block with probability $(\theta+\alpha K)/(\theta+i-1)$, where $K$ is the current number of blocks.
@@ A.2.2 Coagulation of two-parameter random partitions
Recall that a composition is an unordered partition, and a weak composition allows zeros. The following makes Definition 2 in the main text explicit.
@@
Let $\omega,\gamma\in\mathcal{P}$ and $0\leq d\leq\min(l(\omega),l(\gamma))$. For compositions $z,y$ of length $d$ with $a(z)\leq a(\omega)$ and $a(y)\leq a(\gamma)$, with $a(\cdot)$ as in (27), define
@@
$\mathrm{coag}(\omega,\gamma):=\bigcup_{d=0}^{\min(l(\omega),l(\gamma))}\ \bigcup_{z,y}\ \mathrm{coag}_{z,y}(\omega,\gamma),$
@@
$\mathcal{H}(\omega,\gamma|\mu)=\binom{n+m}{n,m}^{-1}\sum\prod_{j=1}^{l(\mu)}\binom{\mu_{j}}{z_{j},y_{j}}$, (31)
@@
$\Pr(\Pi_{1:n+m}=\mu,\ \Pi_{n+1:n+m}=\gamma\mid\Pi_{1:n}=\omega)=\mathcal{H}(\omega,\gamma|\mu)\,\frac{\mathrm{PSF}_{\alpha,\theta}(\mu)}{\mathrm{PSF}_{\alpha,\theta}(\omega)}.$ (32)
@@
$\mathrm{CRP}_{\omega\uparrow\gamma}:=\Pr(\Pi_{n+1:n+m}=\gamma\mid\Pi_{1:n}=\omega)=\sum_{\mu\in\mathrm{coag}(\omega,\gamma)}\mathcal{H}(\omega,\gamma|\mu)\,\frac{\mathrm{PSF}_{\alpha,\theta}(\mu)}{\mathrm{PSF}_{\alpha,\theta}(\omega)},$ (33)
the conditional CRP kernel. Moreover,
@@
and we will write $\Pi_{1:n+m}|\Pi_{1:n}=\omega,\Pi_{n+1:n+m}=\gamma\sim\mathrm{Coag}_{\alpha,\theta}(\omega,\gamma)$, the coagulation kernel. Finally, conditional on $\Pi_{n+1:n+m}=\pi$, any prior distribution $v_{\omega}$ over $\omega$ is updated as
$\hat{v}_{\omega}\;\propto\;v_{\omega}\,\mathrm{CRP}_{\omega\uparrow\pi},$ (35)
and the joint law of $\Pi_{1:n+m}$ is a mixture of coagulation kernels:
$\Pi_{1:n+m}\mid\Pi_{n+1:n+m}=\pi\;\sim\;\sum_{\omega}\hat{v}_{\omega}\,\mathrm{Coag}_{\alpha,\theta}(\omega,\pi).$ (36)
@@
A.3 A CRP with loss of particles as dual process
@@
$\mathbb{E}\!\left[H(X(t),d)\mid X(0)=x\right]=\mathbb{E}\!\left[H(x,D(t))\mid D(0)=d\right].$ (37)
@@
$p^{\downarrow}_{\lambda,\omega}(t):=\Pr(D(t)=\omega\mid D(0)=\lambda)=\mathcal{H}(\omega\mid\lambda)\,d_{|\lambda|,|\omega|}^{\theta}(t),\quad\omega\in L(\lambda),$ (38)
where $H(x,\omega)=P(\omega\mid x)/\mathrm{PSF}_{\alpha,\theta}(\omega)$, $\mathcal{H}(\omega\mid\lambda)$ is obtained by integrating out $\gamma$ from the symmetrized hypergeometric coefficient in Proposition 4, and $d_{n,m}^{\theta}(t)$ is the transition probability of the block-counting process of Kingman’s coalescent. Thus the dual is a pure-death process on integer partitions: from state $\lambda$, jumps occur at rate $|\lambda|(|\lambda|+\theta-1)/2$, and a uniformly chosen square from the Young diagram of $\lambda$ is deleted. A random partition with law (38) will be denoted $\Pi\sim p^{\downarrow}_{\lambda,\bullet}(t)$.
@@
A.4 Details and proofs on filtering and smoothing
A.4.1 Online estimation: filtering and prediction
@@
The second operator, denoted $\phi_{\pi}$, acts as a Bayesian update: if $X(t_{k})\sim\nu$, then $\phi_{\pi}(\nu)$ is the law of $X(t_{k})\mid\Pi^{k}=\pi$.
@@
The predictive law $\nu_{k|0:k-1}$ of $X(t_{k})\mid\Pi^{0:k-1}$ belongs to $\mathcal{F}$:
$X(t_{k})\mid\lambda(t_{k})\sim\mathrm{PD}_{\alpha,\theta}^{\lambda(t_{k})},\qquad\lambda(t_{k})\sim\sum_{\lambda\in\Lambda_{0:k-1}}w_{\lambda}\,p^{\downarrow}_{\lambda,\bullet}(\Delta_{k}),$ (42)
@@
    draw $\lambda(t_{k-1})\sim\{w_{\lambda}\}$,
@@
An important corollary is convergence to stationarity: for $t\to\infty$, $\varphi_{t}(\nu_{N|0:N})\to\mathrm{PD}_{\alpha,\theta}$ in total variation.
@@
Let $\varphi_{t}(\nu_{N|0:N})(\cdot):=\mathrm{Pr}_{\alpha,\theta}(X(t_{N}+t)\in\cdot|\Pi^{0:N})$. Then
@@
$X(t_{k})\mid\lambda(t_{k})\sim\mathrm{PD}_{\alpha,\theta}^{\lambda(t_{k})},\qquad\lambda(t_{k})\mid\pi^{k}\sim\sum_{\omega\in L(\Lambda_{0:k-1})}\hat{v}_{\omega}\mathrm{Coag}_{\alpha,\theta}(\omega,\pi^{k}),$ (44)
with $\hat{v}_{\omega}\propto v_{\omega}\mathrm{CRP}_{\omega\uparrow\pi^{k}}$, where $\mathrm{CRP}_{\omega\uparrow\pi^{k}}$ is in (7) and $\mathrm{Coag}_{\alpha,\theta}$ as in (34). Equivalently,
$\nu_{k|0:k}=\sum_{\lambda\in\Lambda_{0:k}}w_{\lambda}\mathrm{PD}_{\alpha,\theta}^{\lambda},\qquad w_{\lambda}\propto\sum_{\omega\in L(\Lambda_{0:k-1})}\hat{v}_{\omega}\,\Pr(\lambda\mid\omega,\pi^{k}),$ (45)
with $\Lambda_{0:k}=\mathrm{coag}(L(\Lambda_{0:k-1}),\pi^{k})$.
The result follows because $P(\pi^{k}|x)$ can be factorized kernel by kernel in $\nu_{k|0:k-1}$, and Proposition 4 expresses the resulting product as a mixture over coagulations.
@@
A.4.2 Offline estimation: smoothing and data forecasting

### v2

@@ 5 Concluding remarks
We developed an exact and tractable inferential framework for a nonparametric HMM where the latent process is a two-parameter Poisson–Dirichlet diffusion and the observations are unlabelled partitions collected at discrete times. By exploiting Markov duality and combinatorial properties of partition coagulations, we derived recursive updates for both posterior and predictive distributions. These results enabled exact filtering, smoothing, and forecasting algorithms, as well as marginal likelihood-based parameter estimation.
Our approach demonstrates practical scalability and statistical fidelity in modeling time-varying unlabelled partition data. Unlike conventional particle filters, which require simulating high-dimensional latent dynamics and evaluating intractable likelihoods, our method sidesteps both difficulties through structural duality and quasi-conjugacy. The resulting algorithms are substantially faster and more stable, yielding coherent posterior trajectories and sharper uncertainty quantification. As shown through synthetic and real-world experiments, the proposed methodology enables tractable, interpretable, and provably exact inference in settings where the latent structure is infinite-dimensional and only partially observed. These properties make it particularly well suited for applications with computational constraints or unobserved individual-level labels, such as social or genetic data.
Appendix A Supplementary Material
@@ A.1 Two-parameter Poisson–Dirichlet diffusions
with domain given by the subalgebra of the space $C(\overline{\nabla})$ of continuous function on $\overline{\nabla}$ generated by $\{\varphi_{m}(x),m\in\mathbb{N}\}$, with $\varphi_{1}(x)\equiv 1$ and $\varphi_{m}(x)=\sum_{i\geq 1}x_{i}^{m}$ for $m\geq 2$. See, e.g., feng2010poisson, Section 5.5. The transition function of the diffusion was found in feng2011functional through a spectral expansion, and admits the representation as infinite mixture of conditional two-parameter Poisson–Dirichlet distributions
@@
where $a_{(b)}=a(a+1)\cdots(a+b-1)$ is the rising factorial or Pochhammer symbol, with $a_{(0)}:=1$. These $d_{n}(t)$ are the transition probabilities of a death process on $\mathbb{Z}_{+}=\{0,1,\dots\}$ that has entrance boundary at $\infty$ and infinitesimal transition rates $n(\theta+n-1)/2$ from $n$ to n1n-1. This process is in fact the block-counting process associated to Kingman’s coalescent with mutation (tavare1984line).
@@
at all $t>0$, so that $\overline{\nabla}\setminus\nabla$ acts as an entrance boundary. Then, if $X(0)\in\nabla$ almost surely, we can take $\nabla$ as the process state space. For this reason in Section 3.1 we assumed $X(0)\sim\mathrm{PD}_{\alpha,\theta}$, which implies $X(0)\in\nabla$ with probability one (pitman2006combinatorial). All integrals may be taken over $\nabla$, since $\overline{\nabla}\setminus\nabla$ is an entrance boundary.
@@ A.2.1 Two-parameter Poisson–Dirichlet partition structures
An integer partition of $n\in\mathbb{N}$ is a nonincreasing vector $\pi=(\pi_{1},\ldots,\pi_{\ell})$ of positive integers $\pi_{i}\geq\pi_{i+1}>0$ with $|\pi|:=\sum_{i=1}^{\ell}\pi_{i}=n$, where the length $\ell$ of $\pi$ counts its parts. See Figure 2 for examples of Young diagram representation of partitions. Let $\mathcal{P}_{n}$ be the set of partitions of $n$ and $\mathcal{P}:=\bigcup_{n\geq 1}\mathcal{P}_{n}$. For $\pi\in\mathcal{P}$, write $a(\pi):=(a_{j}(\pi),\,j\geq 1)$ where
$a_{j}(\pi):=\sum_{i=1}^{\ell}\mathbbm{1}\{\pi_{i}=j\}$ (28)
@@
A partition structure (kingman_representation_1978) is a family $\{P_{n}:n\in\mathbb{N}\}$ with $P_{n}$ a probability on $\mathcal{P}_{n}$ such that $P_{n}$ is the marginal obtained by removing uniformly at random one square from the Young diagram of $\Pi\sim P_{n+1}$. A fundamental result by kingman_random_1975; kingman_representation_1978; kingman_random_1978 says that any partition structure can be represented as a mixture of paintbox processes,
$P(\pi)=\int_{\overline{\nabla}}P(\pi|x)\mathcal{M}(dx)$ (29)
for some unique probability measure $\mathcal{M}$ on $\overline{\nabla}$, where $P(\pi|x)$ is the law of the paintbox from $x$ given in Eq. (25), as well as in the main text in Eq. (6).
@@
$C(\pi)\;=\;\binom{|\pi|}{\pi_{1},\ldots,\pi_{\ell}}\;\frac{1}{a_{1}(\pi)!\cdots a_{|\pi|}(\pi)!}.$ (30)
It has been shown that the representing measure of the Ewens-Pitman Sampling Formula is the two parameter Poisson–Dirichlet distribution (which is supported on $\nabla$), so that Kingman’s representation reads
$\mathrm{PSF}_{\alpha,\theta}(\pi)=\int_{\nabla}P(\pi|x)\mathrm{PD}_{\alpha,\theta}({\rm d}x);$ (31)
see pitman1995exchangeable; pitman2006combinatorial.
A random partition $\Pi$ of size $n$ with probability mass function $\{\mathrm{PSF}_{\alpha,\theta}(\pi)$, $\pi\in\mathcal{P}_{n}\}$, is denoted by $\Pi\sim\mathrm{PSF}_{\alpha,\theta}$. The dependence on $n$ is omitted since $\mathrm{PSF}_{\alpha,\theta}$ defines a partition structure and is therefore consistent across different values of $n$. Equivalently, $\Pi$ can be generated by the $(\alpha,\theta)$ Chinese restaurant process (CRP): the $i$th item joins an existing block of current size $\pi_{j}$ with probability $(\pi_{j}-\alpha)/(\theta+i-1)$, or starts a new block with probability $(\theta+\ell\alpha)/(\theta+i-1)$, where $\ell$ is the current number of blocks.
@@ A.2.2 Coagulation of two-parameter random partitions
Recall that a composition is a partition where the condition $\pi_{i}\geq\pi_{i+1}$ is not enforced, and a weak composition allows zeros. The following makes Definition 2 in the main text explicit.
@@
Let $\omega,\gamma\in\mathcal{P}$ and $0\leq d\leq\min(\ell(\omega),\ell(\gamma))$. For compositions $z,y$ of length $d$ with $a(z)\leq a(\omega)$ and $a(y)\leq a(\gamma)$, with $a(\cdot)$ as in (28), define
@@
$\mathrm{coag}(\omega,\gamma):=\bigcup_{d=0}^{\min(\ell(\omega),\ell(\gamma))}\ \bigcup_{z,y}\ \mathrm{coag}_{z,y}(\omega,\gamma),$
@@
$\mathcal{H}(\omega,\gamma|\mu)=\binom{n+m}{n,m}^{-1}\sum\prod_{j=1}^{\ell(\mu)}\binom{\mu_{j}}{z_{j},y_{j}}$, (34)
@@
$\Pr(\Pi_{1:n+m}=\mu,\ \Pi_{n+1:n+m}=\gamma|\Pi_{1:n}=\omega)=\mathcal{H}(\omega,\gamma|\mu)\,\frac{\mathrm{PSF}_{\alpha,\theta}(\mu)}{\mathrm{PSF}_{\alpha,\theta}(\omega)}.$ (35)
@@
$\mathrm{PSF}_{\alpha,\theta}^{\omega}(\gamma):=\Pr(\Pi_{n+1:n+m}=\gamma|\Pi_{1:n}=\omega)=\sum_{\mu\in\mathrm{coag}(\omega,\gamma)}\mathcal{H}(\omega,\gamma|\mu)\,\frac{\mathrm{PSF}_{\alpha,\theta}(\mu)}{\mathrm{PSF}_{\alpha,\theta}(\omega)},$ (36)
the conditional Pitman Sampling Formula. Moreover,
@@
and we will write $\Pi_{1:n+m}|\Pi_{1:n}=\omega,\Pi_{n+1:n+m}=\gamma\sim\mathrm{Coag}_{\alpha,\theta}(\omega,\gamma)$, the coagulation kernel. Consider now $\Pi_{1:n}\sim v_{\omega}$ where $\{v_{\omega};\omega\in\Omega\subset\mathcal{P}\}$ is any probability mass function on a finite subset of $\mathcal{P}$. Assume that $\Pi_{n+1:n+m}|\Pi_{1:n}=\omega\sim\mathrm{PSF}_{\alpha,\theta}^{\omega}$, i.e., given the current configuration $\omega$ of the restaurant the next $m$ customers join $\omega$ according to a Chinese Restaurant Process. Then, conditionally on $\Pi_{n+1:n+m}=\pi$, the distribution of $\Pi_{1:n}$ is updated as
$\hat{v}_{\omega}\;\propto\;v_{\omega}\,\mathrm{PSF}_{\alpha,\theta}^{\omega}(\pi),$ (38)
and the joint law of $\Pi_{1:n+m}$ (the full configuration of the restaurant) is a mixture of coagulation kernels:
$\Pi_{1:n+m}|\Pi_{n+1:n+m}=\pi\;\sim\;\sum_{\omega}\hat{v}_{\omega}\,\mathrm{Coag}_{\alpha,\theta}(\omega,\pi).$ (39)
@@
A.2.3 A CRP with loss of particles as dual process
@@
$\mathbb{E}\!\left[H(X(t),d)|X(0)=x\right]=\mathbb{E}\!\left[H(x,D(t))|D(0)=d\right].$ (40)
@@
$p^{\downarrow}_{\lambda,\omega}(t):=\Pr(D(t)=\omega|D(0)=\lambda)=\mathcal{H}(\omega|\lambda)\,d_{|\lambda|,|\omega|}^{\theta}(t),\quad\omega\in L(\lambda),$ (41)
where $H(x,\omega)=P(\omega|x)/\mathrm{PSF}_{\alpha,\theta}(\omega)$, $\mathcal{H}(\omega|\lambda):=\sum_{\gamma}\mathcal{H}(\omega,\gamma|\lambda)$ with $\mathcal{H}(\omega,\gamma|\lambda)$ as in Proposition 4, and $d_{n,m}^{\theta}(t):=\Pr(D(t)=m|D(0)=n)$ is the transition probability of the block-counting process of Kingman’s coalescent. Thus the dual is a pure-death process on integer partitions: from state $\lambda$, jumps occur at rate $|\lambda|(|\lambda|+\theta-1)/2$, and a uniformly chosen square from the Young diagram of $\lambda$ is deleted. A random partition with law (41) will be denoted $\Pi\sim p^{\downarrow}_{\lambda,\bullet}(t)$.
@@
A.3 Details and proofs on filtering and smoothing
A.3.1 Online estimation: filtering and prediction
@@
The second operator, denoted $\phi_{\pi}$, acts as a Bayesian update: if $X(t_{k})\sim\nu$, then $\phi_{\pi}(\nu)$ is the law of $X(t_{k})|\Pi^{k}=\pi$.
@@
The predictive law $\nu_{k|0:k-1}$ of $X(t_{k})|\Pi^{0:k-1}$ belongs to $\mathcal{F}$:
$X(t_{k})|\lambda(t_{k})\sim\mathrm{PD}_{\alpha,\theta}^{\lambda(t_{k})},\qquad\lambda(t_{k})\sim\sum_{\lambda\in\Lambda_{0:k-1}}w_{\lambda}\,p^{\downarrow}_{\lambda,\bullet}(\Delta_{k}),$ (45)
@@
    draw $\lambda(t_{k-1})\sim\{w_{\lambda};\lambda\in\Lambda_{0:k-1}\}$,
@@
The following corollary establishes the convergence of the propagated law to stationarity.
@@
Let $\varphi_{t}(\nu_{N|0:N})(\cdot):=\Pr(X(t_{N}+t)\in\cdot|\Pi^{0:N})$. Then
@@
$X(t_{k})|\lambda(t_{k})\sim\mathrm{PD}_{\alpha,\theta}^{\lambda(t_{k})},\qquad\lambda(t_{k})|\pi^{k}\sim\sum_{\omega\in L(\Lambda_{0:k-1})}\hat{v}_{\omega}\mathrm{Coag}_{\alpha,\theta}(\omega,\pi^{k}),$ (47)
with $\hat{v}_{\omega}\propto v_{\omega}\mathrm{PSF}_{\alpha,\theta}^{\omega}(\pi^{k})$, where $\mathrm{PSF}_{\alpha,\theta}^{\omega}$ is as in (4) and $\mathrm{Coag}_{\alpha,\theta}$ as in (37). Equivalently,
$\nu_{k|0:k}=\sum_{\lambda\in\Lambda_{0:k}}w_{\lambda}\mathrm{PD}_{\alpha,\theta}^{\lambda},\qquad w_{\lambda}\propto\sum_{\omega\in L(\Lambda_{0:k-1})}\hat{v}_{\omega}\,\Pr(\lambda|\omega,\pi^{k}),$ (48)
with $\Lambda_{0:k}=\mathrm{coag}(L(\Lambda_{0:k-1}),\pi^{k})$, and $\Pr(\lambda|\omega,\pi^{k})$ is shorthand for the distribution on $\mathrm{coag}(\omega,\pi^{k})$ as in Eq. (37).
The result follows because $P(\pi^{k}|x)$ can be factorized kernel-by-kernel in $\nu_{k|0:k-1}$, and Proposition 4 expresses the resulting product as a mixture over coagulations.
@@
A.3.2 Offline estimation: smoothing and data forecasting